OpenStudy (anonymous):

Find the maclaurin series for f(x)=(x^2+4)e^(2x) and use this to calculate the 1000th derivative for f(x) at x=0.

OpenStudy (dan815):

oh gosh not this stuff again

OpenStudy (anonymous):

I already found the Maclaurin series and I got: \[\sum_{n=0}^{\infty}\frac{ (x^2+4)(2x)^n }{ n! }\]

OpenStudy (anonymous):

Now how do I compute the 1000th derivative?

OpenStudy (dan815):

just plug 1000?

OpenStudy (anonymous):

Because you obviously don't differentiate that 100 times :P .

OpenStudy (anonymous):

Now that, I am not sure about.

OpenStudy (anonymous):

Because that's a summation. It's not the derivative.

OpenStudy (anonymous):

Ohh wait... It's a maclaurin series right? SO I already took the derivative.

OpenStudy (anonymous):

Ohh... What a dumb question...

OpenStudy (anonymous):

Thanks lol.

OpenStudy (anonymous):

But, isn't that 0?

OpenStudy (anonymous):

because we have : \[\sum_{n=0}^{\infty}\frac{ (x^2+4)(2x)^n }{ n! }\] So if we plug in 0 we get 0 for everything.

OpenStudy (dan815):

oh no wait

OpenStudy (dan815):

u know that f^1000(0) / n! = the coefficient of the term containing x^1000 term in the power series

OpenStudy (anonymous):

Did you mean f(1000)/n! the coefficient of the term containing x^1000 term in the power series?

OpenStudy (dan815):

sithandgiggles is here he is pro

OpenStudy (anonymous):

That he is my friend.

OpenStudy (anonymous):

I don't think that's the right Mclaurin series for your function... I've already answered a similar question asking for the Mclaurin series of that same function. I spent some time looking for a pattern for the n-th derivative, and it doesn't look anything like what your series suggests. Give me a sec to look for it.

OpenStudy (anonymous):

Well here is another one I found: \[\sum_{k=0}^{\infty}\frac{ (2^k)(x^k+2) }{ k!}+\sum_{k=0}^{\infty}\frac{ 2^{k+2}x^k }{ k!}\]

OpenStudy (dan815):